# Confidential Prompting: Protecting User Prompts from Cloud LLM Providers

### v2

@@ 8 Concluding Remarks
When OSMD is impractical.  In scenarios with only a few users but large workloads, the approach of serving entire LLM instances within per-user CVMs may be more practical. This is often the case for institutions or companies, a use case supported by several providers, such as OpenAI’s Foundry. In contrast, SMD enables efficient confidential prompting for a vast number of end users, making it ideal for services that directly interact with large audiences, such as chatbots or personal assistants.
@@
Integrity checks in SMD.  One limitation of SMD is that it assumes a passive adversary model. This is because Ted does not check the integrity of the $Q$ it receives. For stronger threat models that do not assume that adversaries will follow the protocol, integrity checking mechanisms on the received $Q$ are necessary. For example, Ted can partially retain the LLM weights in the TEE and use randomized algorithms, such as Freivalds’ algorithm, to check the integrity of the received $Q$ without fully recomputing it, or implement a zero-knowledge protocol (ZKP).
Automatic / standalone PO.  Prompt obfuscation requires users to annotate the <redacted/> sequences. This process can be automated using a PII screening tool or a smaller local language model to identify PII. Furthermore, prompt obfuscation has the potential to be used as a standalone tool for privacy-preserving LLM inference. For instance, users can create virtual prompts using a local LLM and GQS algorithm, and then utilize them for black-box cloud LLMs (e.g., ChatGPT) to generate text, without revealing identifiable information to the cloud provider.
Acknowledgments
This work is supported in part by NSF Athena AI Institute (Award #2112562) and Yale University.

### v3

@@ A.4 Limitation of OSPD
When OSPD is impractical.  In scenarios with only a few users but large workloads, the approach of serving entire LLM instances with per-user CVMs may be more practical. It is often the case for institutions or companies, a use case supported by several providers, such as OpenAI’s Foundry. In contrast, SPD enables efficient confidential prompting for a vast number of end users, making it ideal for services that directly interact with large audiences, such as chatbots or personal assistants.
@@
Integrity checks in SPD.  One limitation of SPD is it assumes an HBC model because the CVM does not check the integrity of the received $Q$. For stronger threat models that do not assume adversaries will follow the protocol, integrity checking mechanisms on the received $Q$ are necessary. For example, the CVM can partially retain the LLM weights in the TEE and use randomized algorithms, such as Freivalds’ algorithm, to check the integrity of the received $Q$ without fully recomputing it, or implement a zero-knowledge protocol (ZKP).